# Time-To-Inconsistency: A Survival Analysis of Large Language Model Robustness to Adversarial Attacks

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized conversational AI, yet their robustness in extended multi-turn dialogues remains poorly understood. Existing evaluation frameworks focus on static benchmarks and single-turn assessments, failing to capture the temporal dynamics of conversational degradation that characterize real-world interactions. In this work, we present the first comprehensive survival analysis of conversational AI robustness, analyzing 36,951 conversation turns across 9 state-of-the-art LLMs to model failure as a time-to-event process. Our survival modeling framework—employing Cox proportional hazards, Accelerated Failure Time, and Random Survival Forest approaches—reveals extraordinary temporal dynamics. We find that abrupt, prompt-to-prompt(P2P) semantic drift is catastrophic, dramatically increasing the hazard of conversational failure. In stark contrast, gradual, cumulative drift is highly protective, vastly reducing the failure hazard and enabling significantly longer dialogues. AFT models with interactions demonstrate superior performance, achieving excellent discrimination and exceptional calibration. These findings establish survival analysis as a powerful paradigm for evaluating LLM robustness, offer concrete insights for designing resilient conversational agents, and challenge prevailing assumptions about the necessity of semantic consistency in conversational AI Systems.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023), yet their deployment in high-stakes applications necessitates rigorous evaluation of their consistency under adversarial conditions Hendrycks et al. (2021); Lin et al. (2022). While existing evaluation frameworks primarily assess single-turn performance Liang et al. (2022); Gao et al. (2023), real-world interactions involve sustained multi-turn conversations where models must maintain consistency despite evolving contexts and adversarial pressure Shuster et al. (2022); Bai et al. (2022).
Current evaluation paradigms exhibit fundamental limitations in capturing the temporal dynamics of conversational AI robustness Kiela et al. (2021); Ribeiro et al. (2020). Standard benchmarks measure performance in isolated turns, inadequately capturing cumulative effects of conversational drift and emergent vulnerabilities during extended interactions Zheng et al. (2023); Dubois et al. (2023). Phenomena such as sycophancy—wherein models readily abandon correct responses under minimal user challenges Sharma et al. (2023); Turpin et al. (2023)—exemplify systematic fragilities that single-turn evaluations fail to detect.
Consider a medical AI assistant that initially provides accurate information but gradually shifts recommendations under persistent questioning Singhal et al. (2023); Nori et al. (2023), or a system that maintains precision for straightforward queries yet fails catastrophically when confronted with specific combinations of semantic drift and adversarial strategies Zou et al. (2023); Wei et al. (2023). Such failure modes represent critical security concerns for deployed systems Ganguli et al. (2022); Perez et al. (2022), yet remain largely invisible to conventional evaluation metrics Morris et al. (2023); Wang et al. (2023).
To address these gaps, we reframe the problem of multi-turn LLM consistency using survival analysis, a powerful statistical methodology for modeling time-to-event processes Cox (1972); Kalbfleisch & Prentice (2011). We conceptualize conversational failure as the ”event” of interest, with ”time” measured in sequential dialogue turns. This novel perspective allows us to move beyond static metrics and model the temporal dynamics of robustness. Our work represents the first systematic application of survival analysis to LLM conversation robustness evaluation. Unlike existing approaches that rely on static metrics or single-turn evaluation, our framework provides temporal understanding of failure patterns through established survival analysis methodologies. The multi-paradigm approach (semi-parametric Cox models, parametric AFT models, and non-parametric Random Survival Forests) ensures robust conclusions while accommodating different assumptions about failure time distributions and hazard functions. The integration of semantic drift features with survival modeling addresses a critical gap in current LLM evaluation methodologies, providing actionable insights for developing more robust conversational AI systems.
@@ 2.1 Multi-Turn Degradation and Evaluation in LLMs
Recent research consistently demonstrates that large language models (LLMs) exhibit significant performance degradation during multi-turn interactions compared to single-turn tasks Laban et al. (2025); Li et al. (2025b). This degradation manifests primarily as increased inconsistency and variance across conversational turns, arising from premature conclusions and overly confident reliance on incorrect intermediate responses Laban et al. (2025). To systematically measure such inconsistencies, several specialized benchmarks have been developed. Early frameworks such as MT-Bench Zheng et al. (2023) primarily evaluated two-turn interactions, while subsequent efforts like MT-Bench-101 Bai et al. (2024) extended these evaluations to more extensive dialogue scenarios, highlighting uneven multi-turn performance even in advanced chat-tuned models. Complementarily, MT-Eval Kwan et al. (2024) introduced controlled experiments to explicitly contrast single-turn and multi-turn performance, identifying error propagation and distant contextual dependencies as critical contributors to performance decline. Additionally, benchmarks like MultiChallenge Sirdeshmukh et al. (2025) emphasize realistic conversational complexities, exposing significant limitations in current models’ ability to manage ambiguous instructions and context shifts across turns.
@@ 2.2 Consistency and Sycophantic Behavior
Focused examinations into specific multi-turn failure modes have uncovered critical phenomena such as ”sycophantic drift,” where models alter correct answers in response to user pushback or misleading follow-ups. The FlipFlop Experiment by Laban et al. (2023) empirically demonstrated this vulnerability, observing frequent reversals from correct to incorrect answers under trivial user challenges. To quantify and mitigate this issue, Li et al. (2025a) introduced the Position-Weighted Consistency (PWC) metric, penalizing early-stage inconsistencies due to their detrimental impact on user trust. Their Confidence-Aware Response Generation (CARG) method notably improved multi-turn consistency by leveraging the model’s internal confidence signals. Our hazard-modeling approach complements these findings by statistically characterizing the increasing risk of response inconsistency over dialogue turns.
@@ 2.3 Survival Analysis and Sequential Modeling
Survival analysis techniques, traditionally employed to model time-to-event data, provide a natural analytical framework for evaluating sequential behavior in LLMs. De Kock & Vlachos (2021) demonstrated the utility of survival models in conversational AI contexts by predicting dialogue termination and disruptions with greater interpretability than traditional classifiers. Similarly, Maystre & Russo (2022) integrated temporal consistency conditions into survival analysis, significantly improving predictions in sequential decision-making environments. Despite these advances, applying survival analysis explicitly to model turn-by-turn failure risks in LLMs remains largely unexplored. Our work addresses this gap by framing LLM multi-turn consistency as a survival problem, enabling a nuanced statistical characterization of error accumulation and offering novel insights into dialogue reliability dynamics previously observed only empirically.
@@ 3.1 Problem Formulation
We cast conversational robustness as a survival analysis problem in which a failure occurs when the model first produces an incorrect answer during a multi-turn exchange. Time is measured in discrete conversation rounds.
For each conversation $k$, we define:
  • Time-to-event $T_{k}\in\{1,\ldots,H\}$: the number of rounds until the first incorrect answer, with a fixed observation horizon H=8H{=}8.
  • Failure indicator $\delta_{k}\in\{0,1\}$: $\delta_{k}{=}1$ if failure occurs within the horizon ($T_{k}\leq H$), and $\delta_{k}{=}0$ if no error is observed by round $H$ (right-censoring).
  • Covariates $X_{k,t}$: a vector of features that summarize conversation dynamics up to and including round $t$ (e.g., semantic drift patterns, model characteristics, and turn-level interaction features).
Let $S_{k}(t)=\Pr(T_{k}>t\mid X_{k,\leq t})$ denote the (conditional) survival function, i.e., the probability that conversation $k$ remains error-free beyond round $t$. Because time is discrete, we use the discrete-time hazard
$h_{k}(t)\;=\;\Pr(T_{k}=t\mid T_{k}\geq t,\;X_{k,\leq t}),$
which quantifies the instantaneous risk of failure at round $t$ given survival up to $t$. The survival and hazard are linked by
$S_{k}(t)\;=\;\prod_{u=1}^{t}\bigl(1-h_{k}(u)\bigr).$
Our objective is to learn how covariates $X_{k,t}$ relate to time-to-failure $T_{k}$ by estimating $h_{k}(t)$ (or equivalently $S_{k}(t)$), thereby enabling (i) prediction of failure risk across turns and (ii) analysis of how semantic drift, model properties, and conversational features shape the survival dynamics of large language model interactions.
@@ 3.2 Predictive Feature Engineering
To capture the conversational dynamics that may predict failure, we engineer a set of predictive features from the dialogue text. We first generate dense vector representations for key conversational elements using a sentence transformer model Reimers & Gurevych (2019). For each turn $t$ in a conversation, we denote the embedding of the user’s prompt as $e_{t}$. We represent the accumulated historical context up to that point as $\overline{\mathbf{e}}_{1:t-1}$, which is the averaged embedding of all previous prompts. From these vector representations, we derive the following features:
Prompt-to-Prompt Drift ($D_{p2p}$): Measures immediate semantic shift between consecutive conversation turns:
$D_{p2p}(t)=1-\cos(\mathbf{e}_{t-1},\mathbf{e}_{t})$ (1)
where $\mathbf{e}_{t}$ represents the sentence embedding of the prompt at round $t$.
Context-to-Prompt Drift ($D_{c2p}$): Captures deviation from the overall conversation context:
$D_{c2p}(t)=1-\cos(\bar{\mathbf{e}}_{1:t-1},\mathbf{e}_{t})$ (2)
where $\bar{\mathbf{e}}_{1:t-1}$ is the averaged embedding of all previous conversation rounds.
Cumulative Drift ($D_{cum}$): Tracks the total semantic distance traveled:
$D_{cum}(t)=\sum_{i=2}^{t}D_{p2p}(i)$ (3)
Discrete covariates and complexity controls.
We augment $X_{i,t}$ with prompt complexity (token count) and fixed effects for model, subject, and difficulty. Let $p_{i,t}\in\mathbb{N}$ denote the token count of the user–model prompt at round $t$. Categorical factors are encoded via one-hot (dummy) variables:
$S_{i}$ $\in\{s_{1},s_{2},\ldots,s_{7}\}$ (subject-domain cluster) (4)
$L_{i}$ $\in\{l_{1},l_{2},l_{3},l_{4}\}$ (initial-question difficulty) (5)
$M_{i}$ $\in\{m_{1},m_{2},\ldots,m_{R}\}$ (model family/type). (6)
Thus, $X_{i,t}$ includes $p_{i,t}$ and the corresponding dummy vectors for $S_{i}$, $L_{i}$, and $M_{i}$.
@@ 3.3 Survival Modeling Framework
We estimate failure risk using a family of survival models that span semi-parametric, parametric, and non-parametric paradigms. Throughout, $i$ indexes conversations, $t\in\{1,\ldots,H\}$ indexes turns, and $\mathbf{X}_{i}(t)$ denotes the time-varying covariate vector described in §3.2, including drift features $(D_{\mathrm{p2p}},D_{\mathrm{c2p}},D_{\mathrm{cum}})$ and complexity $C$, plus one-hot dummies for model type $\mathbf{M}_{i}$, subject cluster $\mathbf{S}_{i}$, and difficulty $\mathbf{L}_{i}$.
@@
Baseline Cox proportional hazards (PH) with frailty.
We fit a semi-parametric Cox model with turn-varying covariates and conversation-level frailty:
$h_{i}\!\left(t\mid\mathbf{X}_{i}(t),\nu_{i}\right)=\nu_{i}\,h_{0}(t)\,\exp\!\big\{\bm{\beta}^{\!\top}\mathbf{X}_{i}(t)\big\},$ (7)
where $h_{0}(t)$ is an unspecified baseline hazard and $\nu_{i}\sim\mathrm{Gamma}(\theta)$ is a multiplicative frailty capturing unobserved heterogeneity at the conversation level. The covariate vector
$\mathbf{X}_{i}(t)=\big[D_{\mathrm{p2p}}^{(i)}(t),\,D_{\mathrm{c2p}}^{(i)}(t),\,D_{\mathrm{cum}}^{(i)}(t),\,C^{(i)}(t),\,\mathbf{M}_{i},\,\mathbf{S}_{i},\,\mathbf{L}_{i}\big]^{\top}$
collects the time-varying drift/complexity features and the categorical indicators. Coefficients $\bm{\beta}$ are estimated by partial likelihood with gamma-frailty penalization; robust (clustered) standard errors are computed at the conversation level. This baseline treats all LLMs as one population while allowing for model-specific intercept shifts (via $\mathbf{M}_{i}$) and conversation-level noise (via $\nu_{i}$).
Advanced Cox PH with model–drift interactions.
To test whether drift affects models differently, we augment the linear predictor with interactions between model indicators and the drift covariates:
$h_{i}\!\left(t\mid\mathbf{X}_{i}(t)\right)$ $=h_{0}(t)\,\exp\!\big\{\eta_{i}(t)\big\},$ (8)
$\eta_{i}(t)\;$ $=\;\underbrace{\bm{\beta}_{D}^{\!\top}\mathbf{D}_{i}(t)}_{\text{drift main effects}}+\underbrace{\bm{\alpha}^{\!\top}\mathbf{M}_{i}}_{\text{model}}+\underbrace{\bm{\psi}^{\!\top}\mathbf{S}_{i}}_{\text{subject}}+\underbrace{\bm{\lambda}^{\!\top}\mathbf{L}_{i}}_{\text{difficulty}}+\sum_{m=1}^{M-1}\mathbb{I}\{\text{model}=m\}\,\big(\bm{\gamma}_{m}^{\!\top}\mathbf{D}_{i}(t)\big),$ (9)
where $\mathbf{D}_{i}(t)=(D_{\mathrm{p2p}}^{(i)}(t),D_{\mathrm{c2p}}^{(i)}(t),D_{\mathrm{cum}}^{(i)}(t),C^{(i)}(t))$, $\bm{\beta}_{D}\in\mathbb{R}^{4}$ are average drift effects, and $\bm{\gamma}_{m}\in\mathbb{R}^{4}$ modulate drift effects for each non-reference model $m$. For model $m$, the net drift effect is $(\bm{\beta}_{D}+\bm{\gamma}_{m})$, enabling direct comparison of model-specific sensitivities (e.g., susceptibility to p2p drift). Estimation uses partial likelihood with shrinkage (ridge or group-lasso) on interaction blocks to prevent overfitting given H=8H{=}8.
Parametric Accelerated Failure Time (AFT) models.
To assess robustness under alternative assumptions and potential PH violations, we fit AFT models that regress $\log T$ on covariates,
$\log T_{i}\;=\;\mu_{i}\;+\;\sigma\,\varepsilon_{i},\qquad\mu_{i}\equiv\bm{\beta}^{\!\top}\mathbf{X}_{i}^{\ast},\ \ \sigma>0,$
where the error law of $\varepsilon_{i}$ determines the survival family. Let $\lambda_{i}\equiv\exp(\mu_{i})$ denote a scale parameter and $k\equiv 1/\sigma$ a shape parameter when convenient. The acceleration factor $\exp(\Delta\mu)$ multiplies characteristic times (e.g., medians), giving direct time-scaling interpretations of covariates.
Weibull AFT (extreme-value errors). Here $\varepsilon\sim\text{EV}$ with CDF $F(\varepsilon)=1-\exp\{-\exp(\varepsilon)\}$, implying $T\sim\text{Weibull}(k,\lambda)$ with $k{=}1/\sigma$ and $\lambda{=}\lambda_{i}$.
$S(t\mid\mu,\sigma)$ $=\exp\!\left\{-\left(\tfrac{t}{\lambda}\right)^{k}\right\},$ $h(t\mid\mu,\sigma)$ $=\frac{k}{\lambda}\left(\tfrac{t}{\lambda}\right)^{k-1},$
$f(t\mid\mu,\sigma)$ $=\frac{k}{\lambda}\left(\tfrac{t}{\lambda}\right)^{k-1}\exp\!\left\{-\left(\tfrac{t}{\lambda}\right)^{k}\right\},$ median $=\lambda(\ln 2)^{1/k}.$
When k=1k{=}1 the model reduces to exponential AFT. Larger $k$ concentrates mass at earlier times (increasing hazard).
Log-normal AFT (Gaussian errors). Here $\varepsilon\sim\mathcal{N}(0,1)$, so $T\sim\text{LogNormal}(\mu,\sigma^{2})$.
$S(t\mid\mu,\sigma)$ $=1-\Phi\!\left(\frac{\ln t-\mu}{\sigma}\right),$ $f(t\mid\mu,\sigma)$ $=\frac{1}{t\,\sigma\sqrt{2\pi}}\exp\!\left\{-\frac{(\ln t-\mu)^{2}}{2\sigma^{2}}\right\},$
$h(t\mid\mu,\sigma)$ $=\frac{f(t\mid\mu,\sigma)}{S(t\mid\mu,\sigma)},$ median $=\exp(\mu)=\lambda.$
The log-normal hazard is non-monotone (typically unimodal), which can better capture “wear-in → wear-out” patterns in conversational failure.
Log-logistic AFT (logistic errors). Here $\varepsilon\sim\text{Logistic}(0,1)$ with CDF $\Lambda(z)=1/(1+e^{-z})$, yielding $T\sim\text{LogLogistic}(k,\lambda)$ with $k{=}1/\sigma$, $\lambda{=}\exp(\mu)$.
$S(t\mid\mu,\sigma)$ $=\frac{1}{1+\left(\tfrac{t}{\lambda}\right)^{k}},$ $f(t\mid\mu,\sigma)$ $=\frac{(k/\lambda)\left(\tfrac{t}{\lambda}\right)^{k-1}}{\left[1+\left(\tfrac{t}{\lambda}\right)^{k}\right]^{2}},$
$h(t\mid\mu,\sigma)$ $=\frac{(k/\lambda)\left(\tfrac{t}{\lambda}\right)^{k-1}}{1+\left(\tfrac{t}{\lambda}\right)^{k}},$ median $=\lambda.$
The mean exists only for k>1k{>}1 and the variance for k>2k{>}2, offering heavy-tail flexibility for late failures.
For all three, we maximize the right-censored log-likelihood
$\ell(\bm{\beta},\sigma)=\sum_{i=1}^{n}\Big\{\delta_{i}\log f(T_{i}\mid\mu_{i},\sigma)+(1-\delta_{i})\log S(T_{i}\mid\mu_{i},\sigma)\Big\},$
with gradients computed in the $(\mu,\log\sigma)$ reparameterization for numerical stability.
Random Survival Forests (RSF).
Finally, we employ RSF as a non-parametric ensemble method that accommodates complex interactions and nonlinearities among time-varying features. Each tree is grown on a bootstrap sample; at each split, candidate features are drawn at random and split by a survival impurity measure (log-rank score). For an observation, each terminal node yields a Nelson-Aalen cumulative hazard estimate; the forest aggregates these to produce the ensemble cumulative hazard $\widehat{H}_{i}(t)$ and survival $\widehat{S}_{i}(t)=\exp\left\{-\widehat{H}_{i}(t)\right\}$. Model hyperparameters are tuned via cross-validation to optimize predictive performance.

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized conversational AI, yet their robustness in extended multi-turn dialogues remains poorly understood. Existing evaluation frameworks focus on static benchmarks and single-turn assessments, failing to capture the temporal dynamics of conversational degradation that characterize real-world interactions. In this work, we present a large-scale survival analysis of conversational robustness, modeling failure as a time-to-event process over 36,951 turns from 9 state-of-the-art LLMs on the MT-Consistency benchmark. Our framework combines Cox proportional hazards, Accelerated Failure Time (AFT), and Random Survival Forest models with simple semantic drift features. We find that abrupt prompt-to-prompt semantic drift sharply increases the hazard of inconsistency, whereas cumulative drift is counterintuitively protective, suggesting adaptation in conversations that survive multiple shifts. AFT models with model–drift interactions achieve the best combination of discrimination and calibration, and proportional hazards checks reveal systematic violations for key drift covariates, explaining the limitations of Cox-style modeling in this setting. Finally, we show that a lightweight AFT model can be turned into a turn-level risk monitor that flags most failing conversations several turns before the first inconsistent answer while keeping false alerts modest. These results establish survival analysis as a powerful paradigm for evaluating multi-turn robustness and for designing practical safeguards for conversational AI systems.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse tasks brown2020language; chowdhery2022palm; touvron2023llama, yet their deployment in high-stakes applications necessitates rigorous evaluation of their consistency under adversarial conditions hendrycks2021measuring; lin2022truthfulqa. While existing evaluation frameworks primarily assess single-turn performance liang2022holistic; eval-harness, real-world interactions involve sustained multi-turn conversations where models must maintain consistency despite evolving contexts and adversarial pressure shuster2022blenderbot; bai2022training.
Current evaluation paradigms exhibit fundamental limitations in capturing the temporal dynamics of conversational robustness kiela2021dynabench; ribeiro2020beyond. Standard benchmarks measure performance in isolated turns, and even multi-turn protocols are usually summarized by static aggregate scores. These views obscure how errors emerge and propagate over time: they cannot distinguish between a model that fails immediately under mild adversarial pressure and one that remains stable for many turns before eventually degrading. Phenomena such as sycophancy—where models readily abandon correct responses under minimal user challenges sharma2023towards; turpin2023language—illustrate that the trajectory of a conversation matters just as much as its final outcome.
Consider a medical assistant that initially provides accurate information but gradually shifts recommendations under persistent questioning singhal2022large; nori2023capabilities, or a system that maintains precision for straightforward queries yet fails catastrophically only after a specific pattern of semantic drift and adversarial prompts zou2023universal; wei2024jailbroken. In both cases, what makes these failures concerning is not just that they occur, but when they occur and how they are precipitated by the dialogue history. From a safety and reliability perspective, we need tools that can answer questions such as: How quickly do errors emerge under different adversarial strategies? Which kinds of semantic shifts most sharply increase the risk of failure? And can we identify conversations that are on a “high-risk” trajectory before an error actually occurs?
From static accuracy to time-to-event.
We address these questions by framing multi-turn robustness as a time-to-event problem and analyzing the time-to-inconsistency of an LLM within an adversarial conversation. The event is the first incorrect answer under a strict consistency criterion; time is measured in discrete turns; and conversations that remain correct within an 8-turn horizon are treated as right-censored. Survival analysis cox1972regression; kalbfleisch2011statistical naturally fits this setting: it separates whether a conversation fails from when it fails, handles censored dialogues without ad hoc labels, and provides turn-wise hazard functions that track how failure risk evolves over the dialogue. Because survival models support time-varying covariates, they also let us link evolving conversational signals directly to changes in risk.
We instantiate this framework on the MT-Consistency benchmark li-etal-2025-firm, analyzing 36,951 turns from 9 state-of-the-art LLMs using Cox proportional hazards, Accelerated Failure Time (AFT), and Random Survival Forest (RSF) models to understand which assumptions best capture multi-turn failure dynamics. This work makes three main contributions:
  • Framing. We formalize time-to-inconsistency as a survival analysis problem, providing a temporally-aware view of conversational robustness beyond single-turn and static multi-turn metrics.
  • Drift-aware dynamics. We introduce simple semantic drift signals as time-varying covariates and show that abrupt prompt-to-prompt drift sharply increases hazard, whereas cumulative drift is unexpectedly protective, suggesting adaptation in conversations that survive multiple shifts.
  • Methodology and safeguards. We find that AFT models with model–drift interactions offer the best discrimination and calibration, that key drift features violate proportional hazards assumptions, and that lightweight AFT-based monitors can estimate turn-by-turn risk, pointing toward practical real-time safeguards for multi-turn deployments.
@@ 2.1 Multi-Turn Degradation and Evaluation in LLMs
Recent research consistently demonstrates that large language models (LLMs) exhibit significant performance degradation during multi-turn interactions compared to single-turn tasks laban2025llms; li2025beyond. This degradation manifests primarily as increased inconsistency and variance across conversational turns, arising from premature conclusions and overly confident reliance on incorrect intermediate responses laban2025llms. To systematically measure such inconsistencies, several specialized benchmarks have been developed. Early frameworks such as MT-Bench zheng2023judging primarily evaluated two-turn interactions, while subsequent efforts like MT-Bench-101 bai2024mt extended these evaluations to more extensive dialogue scenarios, highlighting uneven multi-turn performance even in advanced chat-tuned models. Complementarily, MT-Eval kwan2024mt introduced controlled experiments to explicitly contrast single-turn and multi-turn performance, identifying error propagation and distant contextual dependencies as critical contributors to performance decline. Additionally, benchmarks like MultiChallenge sirdeshmukh2025multichallenge emphasize realistic conversational complexities, exposing significant limitations in current models’ ability to manage ambiguous instructions and context shifts across turns.
@@ 2.2 Consistency and Sycophantic Behavior
Focused examinations into specific multi-turn failure modes have uncovered critical phenomena such as ”sycophantic drift,” where models alter correct answers in response to user pushback or misleading follow-ups. The FlipFlop Experiment by laban2023you empirically demonstrated this vulnerability, observing frequent reversals from correct to incorrect answers under trivial user challenges. To quantify and mitigate this issue, li-etal-2025-firm introduced the Position-Weighted Consistency (PWC) metric, penalizing early-stage inconsistencies due to their detrimental impact on user trust. Their Confidence-Aware Response Generation (CARG) method notably improved multi-turn consistency by leveraging the model’s internal confidence signals. Our hazard-modeling approach complements these findings by statistically characterizing the increasing risk of response inconsistency over dialogue turns.
@@ 2.3 Survival Analysis and Sequential Modeling
Survival analysis techniques, traditionally employed to model time-to-event data cox1972regression; kalbfleisch2011statistical, has recently been applied to conversational settings, e.g., to predict dialogue termination or disruptions de-kock-vlachos-2021-survival; maystre2022temporally. These works, however, focus on user- or session-level outcomes and do not address the internal consistency of LLM responses under adversarial pressure. In contrast, we model time-to-inconsistency: the first incorrect answer in a multi-turn adversarial dialogue. We combine Cox proportional hazards, Accelerated Failure Time, and Random Survival Forest models and link their behavior to semantic drift covariates, enabling a nuanced statistical characterization of error accumulation and offering novel insights into dialogue reliability dynamics previously observed only empirically.
@@ 3.1 Problem Formulation
We cast conversational robustness as a time-to-event problem in which an event occurs when the model first produces an incorrect answer during a multi-turn adversarial interaction. Time is measured in discrete conversation rounds.
We work with conversations $i=1,\dots,n$ of maximum length $H=8$ turns, following the MT-Consistency protocol (Section 4.1). Each conversation consists of an initial question and up to $H$ adversarial follow-up prompts, paired with model responses. We only retain conversations whose initial answer is correct, so that the event of interest is whether and when the model is swayed away from that correct answer.
For conversation $i$, we define:
  • Event time $T_{i}\in\{1,\ldots,H\}$: the index of the first round at which the model’s answer is labeled inconsistent with the initial correct answer under the MT-Consistency settings.
  • Event indicator $\delta_{i}\in\{0,1\}$: $\delta_{i}{=}1$ if such an inconsistency occurs within the horizon ($T_{i}\leq H$), and $\delta_{i}{=}0$ if no error is observed by round $H$ (right-censoring).
Let $S_{i}(t)=\Pr(T_{i}>t\mid\mathbf{X}_{i,\leq t})$ denote the conditional survival function, i.e., the probability that conversation $i$ remains error-free beyond round $t$ given its history up to $t$. Because time is discrete, we use the discrete-time hazard
$h_{i}(t)\;=\;\Pr(T_{i}=t\mid T_{i}\geq t,\;\mathbf{X}_{i,\leq t}),$
which quantifies the instantaneous risk of failure at round $t$ given survival up to $t$. Survival and hazard are linked by
$S_{i}(t)\;=\;\prod_{u=1}^{t}\bigl(1-h_{i}(u)\bigr).$
Our objective is to learn how a sequence of covariates $\mathbf{X}_{i,t}$, derived from the dialogue up to turn $t$, relates to the event time $T_{i}$. This enables (i) turn-wise prediction of failure risk under adversarial pressure and (ii) analysis of how conversational patterns—in particular semantic drift, domain, difficulty, and model identity—shape the survival dynamics of multi-turn LLM interactions.
@@ 3.2 Predictive Feature Engineering
For each conversation $i$ with user prompts $u_{i,1},\dots,u_{i,H}$ and model responses $r_{i,1},\dots,r_{i,H}$, we construct time-varying covariates $\mathbf{X}_{i,t}$ from two types of embeddings:
Prompt embeddings.
We encode each user prompt with a sentence-transformer model reimers2019sentencebert:
$\mathbf{e}_{i,t}=f(u_{i,t})\in\mathbb{R}^{d}.$
Context embeddings.
We also encode the full conversational context seen by the model up to and including turn $t$. Concretely, we build a text string by concatenating the initial question and all previous user–model messages, followed by the current user prompt:
$\text{context}_{i,t}=\bigl[u_{i,1},r_{i,1},\dots,u_{i,t-1},r_{i,t-1},u_{i,t},r_{i,t}],$
and obtain a context embedding: $\mathbf{c}_{i,t}=f(\text{context}_{i,t})\in\mathbb{R}^{d}.$
Semantic drift features.
From these embeddings we derive three drift metrics:
  • Prompt-to-prompt drift (direct change between consecutive user prompts)
    $D_{\mathrm{p2p}}(i,t)=\begin{cases}0,&t=1,\\ 1-\cos\bigl(\mathbf{e}_{i,t-1},\mathbf{e}_{i,t}\bigr),&t\geq 2;\end{cases}$
  • Context-to-prompt drift (misalignment between what the model has seen so far and the new user input)
    $D_{\mathrm{c2p}}(i,t)=1-\cos\bigl(\mathbf{c}_{i,t-1},\mathbf{e}_{i,t}\bigr);$
  • Cumulative drift (total distance traveled up to turn $t$)
    $D_{\mathrm{cum}}(i,t)=\sum_{s=2}^{t}D_{\mathrm{p2p}}(i,s),\qquad D_{\mathrm{cum}}(i,1)=0.$
Additional covariates.
We further include simple discrete covariates: prompt length $L_{i,t}$ (token count), subject-domain cluster $S_{i}$ (seven thematic domains), difficulty level $D_{i}$ (four bands), and model identity $M_{i}$ (nine LLMs). Categorical variables are one-hot encoded. At each turn $t$, the covariate vector is
$\mathbf{X}_{i,t}=\bigl[D_{\mathrm{p2p}}(i,t),\,D_{\mathrm{c2p}}(i,t),\,D_{\mathrm{cum}}(i,t),\,L_{i,t},\,S_{i},\,D_{i},\,M_{i}\bigr],$
which serves as input to the survival models in Section 3.3.
@@ 3.3 Survival Modeling Framework
Given the time-varying covariates $\mathbf{X}_{i,t}$ defined in Section 3.2, we estimate the event time $T_{i}$ using three complementary survival-model families: (i) semi-parametric Cox proportional hazards models, (ii) parametric Accelerated Failure Time (AFT) models, and (iii) non-parametric Random Survival Forests (RSF). This allows us to compare different assumptions about how risk evolves over turns and how covariates act on the time-to-inconsistency.
Cox proportional hazards models.
@@
Our baseline model is a Cox proportional hazards (PH) model with time-varying covariates:
$h_{i}(t\mid\mathbf{X}_{i,t})=h_{0}(t)\,\exp\bigl(\bm{\beta}^{\top}\mathbf{X}_{i,t}\bigr),$
where $h_{0}(t)$ is an unspecified baseline hazard and $\bm{\beta}$ encodes the effects of semantic drift, prompt length, subject domain, difficulty, and model identity. We estimate $\bm{\beta}$ via partial likelihood and use cluster-robust standard errors at the conversation level.
To capture model-specific sensitivities to drift, we also fit an advanced Cox model in which the linear predictor includes interactions between drift features and model indicators:
$\eta_{i}(t)=\bm{\beta}^{\top}\mathbf{X}_{i,t}\;+\;\sum_{m}\mathbb{I}\{M_{i}=m\}\,\bm{\gamma}_{m}^{\top}\mathbf{D}_{i,t},$
where $\mathbf{D}_{i,t}=\bigl(D_{\mathrm{p2p}}(i,t),D_{\mathrm{c2p}}(i,t),D_{\mathrm{cum}}(i,t)\bigr)$, $\bm{\beta}$ captures global main effects, and $\bm{\gamma}_{m}$ encodes how drift effects are modified for model $m$. We apply mild $\ell_{2}$ regularization to the interaction blocks to avoid overfitting. Proportional-hazards assumptions are checked using Schoenfeld residual tests.
Accelerated Failure Time (AFT) models.
While Cox PH models assume that covariates act multiplicatively on the hazard, AFT models assume that they act multiplicatively on the time scale. We model
$\log T_{i}=\mu_{i}+\sigma\,\varepsilon_{i},\qquad\mu_{i}=\bm{\theta}^{\top}\mathbf{Z}_{i},$
where $\mathbf{Z}_{i}$ summarizes the covariates for conversation $i$ (including aggregated drift statistics, prompt length, subject, difficulty, and model identity), $\sigma>0$ is a scale parameter, and $\varepsilon_{i}$ follows a distribution that specifies the AFT family. We consider standard choices where closed-form survival functions are available: Weibull, log-normal, and log-logistic AFT models; their corresponding $S(t)$ and $h(t)$ are given in Appendix A. The acceleration factor $\exp(\Delta\mu)$ directly quantifies how covariates stretch or shrink characteristic times (e.g., median time-to-inconsistency).
To allow model-specific sensitivities to drift, we also fit AFT models with drift–model interactions by decomposing the linear predictor as
$\mu_{i}=\bm{\theta}^{\top}\mathbf{Z}_{i}\;+\;\sum_{m}\mathbb{I}\{M_{i}=m\}\,\bm{\phi}_{m}^{\top}\mathbf{Z}_{i}^{\mathrm{drift}},$
where $\mathbf{Z}_{i}^{\mathrm{drift}}$ collects conversation-level summaries of $D_{\mathrm{p2p}}$, $D_{\mathrm{c2p}}$, and $D_{\mathrm{cum}}$, $\bm{\theta}$ encodes the global main effects, and $\bm{\phi}_{m}$ captures how drift effects are modified for model $m$. This allows AFT models to represent that abrupt drift may, for example, compress survival times more strongly for some models than others. Parameters are estimated by maximizing the right-censored log-likelihood.
Random Survival Forests.
As a flexible non-parametric baseline, we employ Random Survival Forests (RSF) ishwaran2008random, which fit an ensemble of survival trees on bootstrap samples. At each split, candidate covariates are sampled at random and chosen to maximize a survival impurity reduction (log-rank statistic). Each terminal node yields a Nelson–Aalen estimate of the cumulative hazard; the forest prediction for conversation $i$ is obtained by averaging cumulative hazards across trees and converting to survival probabilities. RSF can capture nonlinearities and high-order interactions between drift features and model identity without explicit parametric assumptions.